# Uncertainty-Aware Digital Twins: Robust Model Predictive Control using Time-Series Deep Quantile Learning

### v1

@@ 2.3 Quantile Regression
$=TiDE(\mathbf{x}_{k-w+1:k},\mathbf{d}_{k-w:k-1},\mathbf{u}_{k-w:k-1},\mathbf{d}_{k:k+N-1},\mathbf{u}_{k:k+N-1}|\bm{\phi}).$ (9)
@@
$[\bar{\hat{\mathbf{x}}}^{f}_{k+1},\tilde{\hat{\mathbf{x}}}^{f}_{k+1},\underline{\hat{\mathbf{x}}}^{f}_{k+1}]^{T}=TiDE(\mathbf{x}_{k}^{p},\mathbf{u}_{k}^{p},\mathbf{u}_{k}^{f},\mathbf{d}_{k}^{p},\mathbf{d}_{k}^{f}|\bm{\phi}).$ (10)
@@
3 Simultaneous multi-step Robust MPC
The purpose of this work is to integrate the simultaneous multi-step ahead predictive quantile with MPC, demonstrating a learning-based robust MPC that performs decision-making under uncertainty for Digital Twins.
3.1 Uncertainty-aware MPC Formulation
@@
$\min\limits_{\mathbf{u}}\leavevmode\nobreak$ $\mathbb{E}_{\mathbf{w}}[J(\mathbf{u},\mathbf{x},\mathbf{r},\mathbf{w})],$ (11a)
@@
3.2 Reformulation with Quantiles and Constraint Tightening
@@
$s.t.\quad$ $[\bar{\hat{\mathbf{x}}}^{f}_{k},\tilde{\hat{\mathbf{x}}}^{f}_{k},\underline{\hat{\mathbf{x}}}^{f}_{k}]^{T}=TiDE(\mathbf{x}_{k}^{p},\mathbf{u}_{k}^{p},\mathbf{u}_{k}^{f},\mathbf{d}_{k}^{p},\mathbf{d}_{k}^{f}|\bm{\phi})$ (12b)
@@ 3.4.2 Penalty Method: Augmented Lagrangian Method
where $\alpha>1$ is the increasing rate of $\mu_{i}$, and $\mathbf{u}_{s}$ is the solution for solving unconstrained optimization (Equation (13)) at iteration $s$. In the next iteration, the solver will resolve the problem using $\mathbf{u}_{s}$ as the initial guess for warm start. Here, we apply the $\text{ReLU}()$ function in Equation (13) because it is a continuous and differentiable function that only penalizes $\Phi_{s}$ when the constraints $c_{i}$ are violated, and enabling the smooth computation of the gradient of $\Phi_{s}$.
@@ 4 Illustrative Example
We first verify the proposed method using a linear invariant system so that the result can be compared with the tube-based method, which is one of the most widely adopted robust MPC methods. The system with exogenous noise on input $\epsilon_{k}\sim\mathcal{N}(0,0.1^{2})$ is formulated as follows:
$\mathbf{x}_{k+1}$ $=\begin{bmatrix}0.3&0.1\\ 0.1&0.2\\ \end{bmatrix}\mathbf{x}_{k}+\begin{bmatrix}0.5\\ 1\end{bmatrix}\left(u_{k}+\epsilon_{k}\right)$ (16)
$=F_{w}(\mathbf{x}_{k},u_{k},\mathbf{w}_{k}),$ (17)
@@
The goal of this example is to use MPC to perform a reference tracking task on $x_{1}$, while maintaining $x_{1}$, $x_{2}$, and $u$ within the feasible regions subjected to unbounded disturbance. The multi-step robust MPC using the quantile prediction from TiDE can be formulated as:
@@
$s.t.\quad$ $\hat{\mathbf{x}}_{k+1}^{f}=[\bar{\hat{\mathbf{x}}}_{k+1}^{f},\tilde{\hat{\mathbf{x}}}_{k+1}^{f},\underline{\hat{\mathbf{x}}}_{k+1}^{f}]^{T}=TiDE(\mathbf{x}_{k}^{p},\mathbf{u}_{k}^{p},\mathbf{v}),$ (18b)
@@
Both $\mathbf{Q}$ and $\mathbf{R}$ are set to $\mathbf{I}$. The original input bound is $\mathbb{U}\in[-5,5]$, while it is dynamically tightened based on prediction error bound $\mathbb{Z}$. The linear regulator $\mathbf{K}$ = [-0.0621, -0.2027] is assigned in this case, and can be further optimized offline using methods such as Bayesian optimization [47], while optimizing $\mathbf{K}$ is out of the scope of this work. The Further, since the predicted upper and lower quantiles have already taken into account the probabilistic bounds on $\hat{x}_{1,k}$, the constraints in Equations (18d)-(18g) can be rewritten as:
@@
To generate state trajectories for system identification using TiDE, a sequence of input $\mathbf{D}_{u}=\{u_{0},\ldots,u_{n-1}\}$, where $\mathbf{u}\in[-5,5]$ is uniformly sampled with a size of $n=422,000$. By setting the initial state $\mathbf{x}_{0}=[0,0]^{T}$, the trajectory of $\mathbf{D}_{x}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\}$ can be simulated by using $\mathbf{u}$ as the input. Furthermore, both $\mathbf{D}_{u}$ and $\mathbf{D}_{x}$ are divided into fractions using the moving window approach, with each fraction having a length of $w+N$. In this case, the window size is $w=10$, and the horizon length is $N=10$. By denoting the $l$th fraction as $\mathbf{D}^{l}_{u}=\{u_{l},\ldots,u_{l+w+N-1}\}$ and $\mathbf{D}^{l}_{x}=\{\mathbf{x}_{l},\ldots,\mathbf{x}_{l+w+N-1}\}$, respectively, we can assign $\mathbf{x}_{p}=[\mathbf{x}_{l+1},\ldots,\mathbf{x}_{l+w}]$ as the past target, $\mathbf{x}_{f}=[\mathbf{x}_{l+w+1},\ldots,\mathbf{x}_{l+w+N}]$ as the future target, with $\mathbf{u}_{p}=[u_{l},\ldots,u_{l+w-1}]$ and $\mathbf{u}_{f}=[u_{l+w},\ldots,u_{l+w+N-1}]$ as the past and future covariates (inputs), respectively. Then, $\mathbf{D}^{l}_{x}$ and $\mathbf{D}^{l}_{u}$ is further split into training, validation, and test set with a 8:1:1 ratio. Following Equation (6), the TiDE model for identifying this system can be trained via supervised learning using quantile loss, which is formulated as:
$\min_{\bm{\phi}}\quad$ $L_{Q}(\mathbf{x}^{f},\hat{\mathbf{x}}^{f})$ (20a)
$s.t.\quad$ $\hat{\mathbf{x}}^{f}=TiDE(\mathbf{x}^{p},\mathbf{u}^{p},\mathbf{u}^{f}|\bm{\phi}).$ (20b)
Table 1: Hyperparamters of training TiDE model
Details for TiDE model setup
1 1 16 128 32 0.2 True
Details for TiDE model training
0.001 0.002 10 0.95 1500 64 True
In Figure 6(b), we compare the predicted median of $x_{1}$ $\tilde{\hat{x}}_{1,k+1:k+N}$ and the ground truth (obtained from the system model without injected noise) using the optimal control input $\mathbf{u}_{k}^{*}$ obtained by solving Equation (18) with a step function as the reference. Here, we demonstrate that the optimizer successfully solves the MPC problem that minimizes the reference tracking error using the multi-step ahead prediction by TiDE. By examining the results, we observe that the predicted response shares a similar trend with the true response. This test validates the optimization capability of TiDE as the multi-step ahead predictor in MPC.
The performance of the proposed robust MPC method is compared with the multi-step MPC without uncertainty awareness (so-called nominal MPC) in Fig. 7. In this example, the reference trajectory is designed to overlap with the bound of $x_{1}$ and will violate the bound of $x_{2}$ during this reference tracking task, aiming to test the constraint handling capability of robust MPC at extreme scenarios. As can be seen from Fig. 7(a) where the nominal multi-step MPC is solved without considering the uncertainty of the prediction, the result yields significant constraint violation due to the disturbance. Since the gain of ancillary controller is small in this case, the applied input is almost identical to the nominal input. Fig. 7(b)-7(d) compares the nominal TiDE prediction and the ground truth at different instances when the optimal control input sequence is solved. Here, the ground truth (green line) is verified by simulating Equation (17) without disturbance. As a result, even though the discrepancy between the nominal prediction and the ground truth is not significant, indicating the accuracy of the prediction, the optimal control input only provides ideal solutions that minimize the MPC loss but ignore the impact of uncertainty and does not provide safety buffer to accommodate disturbance.
@@ 5 Engineering Case study: Directed Energy Deposition Additive Manufacturing
In this section, we implement the proposed multi-step robust MPC as the online decision-making process for the Digital Twin of the Directed Energy Deposition (DED) AM system. Given the inherent uncertainty associated with material variability and environmental factors in the DED process, proactive control strategies, such as MPC, become crucial to achieve desired material properties while minimizing defects [48]. Furthermore, the intricate dynamics of the melt pool make it challenging to develop a physics-based model capable of providing accurate predictions in real-time. Therefore, data-driven methods become promising tools for addressing this challenge.
5.1 Problem Formulation
$[\hat{\mathbf{x}}_{\text{temp},k+1}^{f},\hat{\mathbf{x}}_{\text{depth},k+1}^{f}]^{T}=TiDE(\mathbf{x}_{\text{temp},k}^{p},\mathbf{x}_{\text{depth},k}^{p},\mathbf{d}_{x,k}^{p:f},{d}_{y,k}^{p:f},\mathbf{z}_{k}^{p:f},\mathbf{u}_{k}^{p:f}),$ (21d)
[…]
5.2 System Setup
In this study, the physical DED is replaced by an in-house developed explicit finite element analysis (FEA) code developed by Liao [51]. The code is accelerated by GPU computation using CuPy. It is employed for part-scale transient heat transfer simulation of the DED process. We select a single-track square as the target geometry, as shown in Fig. 10, and its specifications are listed in Table. 10. This numerical setup allows efficient simulation of complex thermal dynamics while maintaining high accuracy in capturing melt pool behavior. Here, we highlight that since the layer height is 0.75mm, we can set $x_{\text{depth,lb}}$ and $x_{\text{depth,ub}}$ to 0.225 and 0.075, corresponding to 10% and 30% dilution.
@@ 5.2 System Setup
The data for training and validating TiDE is generated offline, using the method proposed by Karkaria [3] to generate laser power trajectories using design of experiment (DOE). A total of 100 simulations with varying laser power profiles are conducted. The melt pool temperature and depth are extracted from the FEA model at each timestep. This diverse dataset ensures that TiDE can learn the relationship between process parameters and thermal responses across a wide operational range. As the data are generated, we train a TiDE model as a multi-variate multi-step ahead predictor with window size $w=50$ and horizon size $N=50$, and the 0.95 and 0.05 quantile are assigned as the upper and lower bounds. The specifics of feature extraction, data processing, and training details, as well as model evaluation, are provided in our previous work [50]. As a result, the MAPE and the RRMSE for melt pool temperature predictions are 1.29% and 0.054, respectively, and those for the melt pool depth are 4.25% and 0.0441.

### v3

@@ 2.4 Implementation of Quantile Regression on TiDE
$=\text{TiDE}(\mathbf{x}_{k-w+1:k},\mathbf{d}_{k-w:k-1},\mathbf{u}_{k-w:k-1},\mathbf{d}_{k:k+N-1},\mathbf{u}_{k:k+N-1}|\bm{\phi}).$ (10)
@@
$[\bar{\hat{\mathbf{x}}}^{f}_{k+1},\tilde{\hat{\mathbf{x}}}^{f}_{k+1},\underline{\hat{\mathbf{x}}}^{f}_{k+1}]^{T}=\text{TiDE}(\mathbf{x}_{k}^{p},\mathbf{u}_{k}^{p},\mathbf{u}_{k}^{f},\mathbf{d}_{k}^{p},\mathbf{d}_{k}^{f}|\bm{\phi}).$ (11)
@@
3 Simultaneous Multi-Step Robust MPC for Digital Twin
The purpose of this work is to integrate the simultaneous multi-step ahead predictive quantile with MPC, demonstrating a learning-based robust MPC that performs decision-making under uncertainty for Digital Twins. In this section, we first describe the development of the virtual system via TiDE. Then, we focus on the robust MPC formulation incoperating multi-step quantile prediction, and introduce the optimization tools that can benefit real-time solving. A complete walk-through of the process will be detailed in Section 4.
3.1 Building the Virtual System via TiDE: Data Generation and Model Training
3.2 Real-Time Decision-Making via MPC
3.2.1 Uncertainty-Aware MPC Formulation
@@
$\min_{\mathbf{u}}\leavevmode\nobreak$ $\mathbb{E}{\mathbf{w}}[J(\mathbf{u},\mathbf{x},\mathbf{r},\mathbf{w})],$ (12a)
@@
3.2.2 Reformulation with Quantiles and Constraint Tightening
@@
$s.t.\quad$ $[\bar{\hat{\mathbf{x}}}^{f}_{k},\tilde{\hat{\mathbf{x}}}^{f}_{k},\underline{\hat{\mathbf{x}}}^{f}_{k}]^{T}=\text{TiDE}(\mathbf{x}_{k}^{p},\mathbf{u}_{k}^{p},\mathbf{u}_{k}^{f},\mathbf{d}_{k}^{p},\mathbf{d}_{k}^{f}|\bm{\phi})$ (13b)
@@
Theoretically, the optimal value of $\mathbf{K}$ can be determined using the LQR for linear tube-based robust MPC [8, 16]. This value plays a crucial role in stabilizing the system and preventing the estimation error bound from diverging during multi-step-ahead prediction. It also influences the level of conservativeness in estimating error bounds when propagating noise through recursive rollout. However, for nonlinear (and even unknown) systems, LQR cannot be used to obtain $\mathbf{K}$, and it must be fine-tuned or optimized using black-box methods such as Bayesian optimization [47]. It’s important to note that optimizing $\mathbf{K}$ is beyond the scope of this work. In contrast, TiDE directly captures the actual distribution from the open-loop data, unlike tube-based robust MPC. Therefore, the value of $\mathbf{K}$ does not affect the conservativeness of the predicted error bounds. In online solving, the value of $\mathbf{K}$ tightens the design space of $\mathbf{u}_{k}$ to account for system disturbances. This is demonstrated in the example provided, which can also be derived from the Riccati equation [16]. A more rigorous implementation would involve continuous linear approximations [48] on the TiDE model and deriving $\mathbf{K}$. However, this is also beyond the scope of our work.
@@ 3.3.2 Penalty Method: Augmented Lagrangian Method
where $\alpha>1$ is the increasing rate of $\mu_{i}$, and $\mathbf{u}_{s}$ is the solution for solving unconstrained optimization (Eq. (14)) at iteration $s$. In this work, we select $\lambda_{0}=10,\mu_{0}=1$, and $\alpha=3$ from trial and errors that balance constraint satisfaction and convergence rate. In the next iteration, the solver will resolve the problem using $\mathbf{u}_{s}$ as the initial guess for warm start. Here, we apply the $\text{ReLU}()$ function in Eq. (14) because it is a continuous and differentiable function that only penalizes $\Phi_{s}$ when the constraints $c_{i}$ are violated, and enabling the smooth computation of the gradient of $\Phi_{s}$.
@@ 4 Illustrative Example
We first verify the proposed method using a linear invariant system so that the result can be compared with the tube-based method, which is one of the most widely adopted robust MPC methods. This section provides a complete walk-through by introducing the physical system, developing the virtual system, building the virtual-to-physical connection via robust MPC, and result validation and comparison.
4.1 Physical System
A discrete linear system is selected for demonstration. The system with exogenous noise on input $\epsilon_{k}\sim\mathcal{N}(0,0.1^{2})$ is formulated as follows:
$\mathbf{x}_{k+1}$ $=\begin{bmatrix}0.3&0.1\\ 0.1&0.2\\ \end{bmatrix}\mathbf{x}_{k}+\begin{bmatrix}0.5\\ 1\end{bmatrix}\left(u_{k}+\epsilon_{k}\right)=F_{w}(\mathbf{x}_{k},u_{k},\mathbf{w}_{k}),$ (17)
@@ 4.1 Physical System
4.2 Virtual System
4.2.1 Data Generation and Model Training
The development of the virtual system in our work follows the steps of Stage 1 in Fig. 1. To generate state trajectories for system identification using TiDE, a sequence of input $\mathbf{D}_{u}=\{u_{0},\ldots,u_{n-1}\}$, where $\mathbf{u}\in[-5,5]$ is uniformly sampled with a size of $n=422,000$. By setting the initial state $\mathbf{x}_{0}=[0,0]^{T}$, the trajectory of $\mathbf{D}_{x}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\}$ can be simulated by using $\mathbf{u}$ as the input. Furthermore, both $\mathbf{D}_{u}$ and $\mathbf{D}_{x}$ are divided into fractions using the moving window approach, with each fraction having a length of $w+N$. In this case, the window size is $w=10$, and the horizon length is $N=10$. By denoting the $l$th fraction as $\mathbf{D}^{l}_{u}=\{u_{l},\ldots,u_{l+w+N-1}\}$ and $\mathbf{D}^{l}_{x}=\{\mathbf{x}_{l},\ldots,\mathbf{x}_{l+w+N-1}\}$, respectively, we can assign $\mathbf{x}_{p}=[\mathbf{x}_{l+1},\ldots,\mathbf{x}_{l+w}]$ as the past target, $\mathbf{x}_{f}=[\mathbf{x}_{l+w+1},\ldots,\mathbf{x}_{l+w+N}]$ as the future target, with $\mathbf{u}_{p}=[u_{l},\ldots,u_{l+w-1}]$ and $\mathbf{u}_{f}=[u_{l+w},\ldots,u_{l+w+N-1}]$ as the past and future covariates (inputs), respectively. Then, $\mathbf{D}^{l}_{x}$ and $\mathbf{D}^{l}_{u}$ is further split into training, validation, and test set with a 8:1:1 ratio. Following Eq. (6), the TiDE model for identifying this system can be trained via supervised learning using quantile loss, which is formulated as:
$\min_{\bm{\phi}}\quad$ $L_{Q}(\mathbf{x}^{f},\hat{\mathbf{x}}^{f})$ (18a)
$s.t.\quad$ $\hat{\mathbf{x}}^{f}=\text{TiDE}(\mathbf{x}^{p},\mathbf{u}^{p},\mathbf{u}^{f}|\bm{\phi}).$ (18b)
Table 2: Hyperparamters of training TiDE model
Details for TiDE model setup
1 1 16 128 32 0.2 True
Details for TiDE model training
0.001 0.002 10 0.95 1500 64 True
4.2.2 Model Evaluation
In Figure 6(b), we compare the predicted median of $x_{1}$ $\tilde{\hat{x}}_{1,k+1:k+N}$ and the ground truth (obtained from the system model without injected noise) using the optimal control input $\mathbf{u}_{k}^{*}$ obtained by solving Eq. (19) with a step function as the reference. Here, we demonstrate that the optimizer successfully solves the MPC problem that minimizes the reference tracking error using the multi-step ahead prediction by TiDE. By examining the results, we observe that the predicted response shares a similar trend with the true response. This test validates the optimization capability of TiDE as the multi-step ahead predictor in MPC.
4.3 Virtual-to-Physical Integration
The bi-directional interaction between the physical and virtual system is constructed with robust MPC, where the current state of the physical system is feedback to the virtual system as its prediction input. Further, an optimal control input is solved by an online robust MPC, using the predicted quantiles as the safety buffer to explicitly handle constraints under uncertainty. Then, the optimal control input is applied to the physical system.
The goal of this example is to use robust MPC to perform a reference tracking task on $x_{1}$, while maintaining $x_{1}$, $x_{2}$, and $u$ within the feasible regions subjected to unbounded disturbance. The multi-step robust MPC using the quantile prediction from TiDE can be formulated as:
@@ 4.3 Virtual-to-Physical Integration
$s.t.\quad$ $\hat{\mathbf{x}}_{k+1}^{f}=[\bar{\hat{\mathbf{x}}}_{k+1}^{f},\tilde{\hat{\mathbf{x}}}_{k+1}^{f},\underline{\hat{\mathbf{x}}}_{k+1}^{f}]^{T}=\text{TiDE}(\mathbf{x}_{k}^{p},\mathbf{u}_{k}^{p},\mathbf{v}),$ (19b)
@@
Both $\mathbf{Q}$ and $\mathbf{R}$ are set to $\mathbf{I}$. The original input bound is $\mathbb{U}\in[-5,5]$, while it is dynamically tightened based on prediction error bound $\mathbb{Z}$. The linear regulator $\mathbf{K}$ = [-0.0621, -0.2027] is assigned in this case. The Further, since the predicted upper and lower quantiles have already taken into account the probabilistic bounds on $\hat{x}_{1,k}$, the constraints in Equations (19d)-(19g) can be rewritten as:
@@
The performance of the proposed robust MPC method is compared with the multi-step MPC without uncertainty awareness (so-called nominal MPC) in Fig. 7. In this example, the reference trajectory is designed to overlap with the bound of $x_{1}$ and will violate the bound of $x_{2}$ during this reference tracking task, aiming to test the constraint handling capability of robust MPC at extreme scenarios. As can be seen from Fig. 7(a) where the nominal multi-step MPC is solved without considering the uncertainty of the prediction, the result yields significant constraint violation due to the disturbance. Since the gain of ancillary controller is small in this case, the applied input is almost identical to the nominal input. Fig. 7(b)-7(d) compares the nominal TiDE prediction and the ground truth at different instances when the optimal control input sequence is solved. Here, the ground truth (green line) is verified by simulating Eq. (17) without disturbance. As a result, even though the discrepancy between the nominal prediction and the ground truth is not significant, indicating the accuracy of the prediction, the optimal control input only provides ideal solutions that minimize the MPC loss but ignore the impact of uncertainty and does not provide safety buffer to accommodate disturbance.
@@ 5 Engineering Case study: Directed Energy Deposition Additive Manufacturing
In this section, we implement the proposed multi-step robust MPC as the online decision-making process for the Digital Twin of the Directed Energy Deposition (DED) AM system. Given the inherent uncertainty associated with material variability and environmental factors in the DED process, proactive control strategies, such as MPC, become crucial to achieve desired material properties while minimizing defects [54]. Furthermore, the intricate dynamics of the melt pool makes it challenging to develop a physics-based model capable of providing accurate predictions in real-time. Therefore, data-driven methods have become promising tools for addressing this challenge.
In this case study, our Digital Twin focuses exclusively on the melt pool rather than the entire part to enable real-time decision-making for process control. The melt pool is the most dynamic and sensitive region of the process, where critical quality indicators such as porosity, microstructure, and residual stress originate. Modeling the entire part with sufficient fidelity for in-situ control would be computationally prohibitive and incompatible with the actionable time requirements of online decision-making. In contrast, a melt pool-level Digital Twin allows for high-resolution, low-latency predictions that align with the frequency and spatial resolution of available in-situ sensing data, such as pyrometer readings. Moreover, control inputs like laser power directly influence melt pool behavior, making it the most actionable domain for decision-making via MPC. By concentrating modeling efforts on the melt pool, the Digital Twin remains both computationally tractable and operationally relevant, enabling accurate state prediction and MPC throughout the build process.
The Digital Twin of the DED process in this work comprises three key modules: (a) the physical system, represented by a high-fidelity finite element analysis (FEA) simulation, (b) the virtual system, implemented as a TiDE model that predicts melt pool temperature and depth, and (c) the virtual-to-physical integration, achieved through the proposed robust MPC framework. At each MPC timestep, the current melt pool temperature and depth are extracted from the physical system and used as feedback to the virtual system. The virtual system then predicts future melt pool behavior conditioned on both the feedback and the candidate control input, which is subsequently optimized by the robust MPC. Once the optimal control input is determined, it is applied to the physical system, completing the bi-directional interaction between the physical and virtual domains.
5.1 Physical System: DED Setup
In this study, the physical DED is replaced by an in-house developed explicit FEA code developed by Liao [55]. The code is accelerated by GPU computation using CuPy. It is employed for part-scale transient heat transfer simulation of the DED process. We select a single-track square as the target geometry, as shown in Fig. 10, and its specifications are listed in Table. 10. This numerical setup allows efficient simulation of complex thermal dynamics while maintaining high accuracy in capturing melt pool behavior. Here, we highlight that since the layer height is 0.75mm, we can set $x_{\text{depth,lb}}$ and $x_{\text{depth,ub}}$ to 0.225 and 0.075, corresponding to 10% and 30% dilution. For the details of the setup of DED, feature extraction, and data processing, please refer to our previous work [56] for the technical details.
@@ 5.1 Physical System: DED Setup
5.2 Virtual System: Surrogate Modeling for the Melt Pool of DED
To effectively generate a variety of laser power profile, the design of experiment (DOE) of the time series of the laser power profile is implemented. This method, proposed by Karkaria [3], represents each time series with a 10-dimensional space using Fourier transform, and generates laser power trajectories using Latin Hypercube Sampling. For details of this approach, please refer to our previous work [3, 56]. A total of 100 simulations with varying laser power profiles are conducted. The melt pool temperature and depth are extracted from the FEA model at each timestep. This diverse dataset ensures that TiDE can learn the relationship between process parameters and thermal responses across a wide operational range. As the data are generated, we train a TiDE model as a multi-variate multi-step ahead predictor with window size $w=50$ and horizon size $N=50$, and the 0.95 and 0.05 quantile are assigned as the upper and lower bounds. The specifics of training details, as well as model evaluation, are provided in our previous work [56]. As a result, the MAPE and the RRMSE for melt pool temperature predictions are 1.29% and 0.054, respectively, and those for the melt pool depth are 4.25% and 0.0441.